zackariahharriett:

What is the value of t in the equation 3(2t + 5) = 5t + 25? (4 points) 5 6 10 35 please help

sophia333:

3 ( 2t + 5 ) = 5 t + 25 3 * 2 t + 3 * 5 = 5 t + 25 6 t + 15 = 5 t + 25 6t - 5t = 25 - 15 t = 10 Hope this helps!

zackariahharriett:

thank you

sophia333:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zackariahharriett thank you \(\color{#0cbb34}{\text{End of Quote}}\) Np :)

Downpour17:

Thank you. The question has been answered, please close this. Thanks! Welcome to questioncove!